Mirabile Corporation uses activity-based costing to compute product margins. Overhead costs have already been allocated to the company's three activity cost pools--Processing, Supervising, and Other. The costs in those activity cost pools appear below:

   Processing$3,740

Supervising$27,840

Other$10,600

Processing costs are assigned to products using machine-hours (MHs) and Supervising costs are assigned to products using the number of batches. The costs in the Other activity cost pool are not assigned to products. Activity data appear below:

MHs
(Processing)
Batches
(Supervising)
Product M0 10,450 600
Product M5 550 600
Total 11,000 1,200

Finally, sales and direct cost data are combined with Processing and Supervising costs to determine product margins.

Product M0 Product M5
Sales (total) $ 75,700 $ 90,900
Direct materials (total) $ 28,700 $ 31,600
Direct labor (total) $ 28,000 $ 41,900

What is the product margin for Product M5 under activity-based costing?

  • $1,527

  • $17,400

  • $19,000

  • $3,293
Answer #1

Calculation of product margin:

Sales = $90,900
Less: Direct material $31,600

Direct Labor = $41,900

Overheads:

Processing = 3740*550/11000 = $187

Supervision = 27840*600/1200 = $13,920

Margin = $3,293

Hence, the answer is $3,293